# Understanding the Drag Torque in Common Envelope Evolution

### v1

@@ 4.7 Comparison of the two idealized models
As we have shown, the lagging spheroid model of E04 and the double perturber model of K08 both reproduce the torque in the simulation reasonably well. Thus, they would be expected to be consistent with one another. Indeed, K08 applied their model to the E04 simulation, arguing that their model can be used to explain the E04 results if $\mathcal{M}_{\mathrm{p}}\sim 0.6$. The bottom-left panel of Figure 3 is adapted from K08 (their figure 1a). The quantity plotted in colour is $\log(\mathcal{D})=\log[c_{\mathrm{s,0}}^{2}a/Gm)\lambda]$, which is equal to a constant plus the logarithm of the fractional density perturbation, $\lambda=(\rho-\rho_{0})/\rho_{0}$ (with $\lambda\ll 1$ assumed in K08), for $\mathcal{M}_{\mathrm{p}}=0.6$. From the bottom panel of Figure 2 the relevant part of our simulation typically has $\mathcal{M}_{\mathrm{p}}\sim 0.6$ as well. To compare the ellipsoid model with that of K08, we annotate their figure by drawing an ellipse with parameters set equal to the time-averaged values from our simulation: $\langle A/a\rangle=1.47$, $\langle B/A\rangle=0.649$, and $\langle\Delta\phi=14.9^{\circ}\rangle$ (see also Figure 3). The fairly close correspondence helps to explain why the E04 and K08 models produce similar values of the torque. The time-averaged best fit ellipse in the orbital plane from our simulation is seen to align quite well with the structure obtained for $\log(\mathcal{D})$ by K08, though the latter is not precisely elliptical as the wakes form elongated tails.
The bottom-right panel of Figure 3 shows $\log(\mathcal{D})$ from our simulation at $t=188.7$ d, which is the same snapshot plotted in the top two panels. To calculate $\mathcal{D}$, we assume $\rho_{0}=0.44\rho_{\mathrm{c}}=0.44\times 0.006\rho_{\mathrm{max}}(t)$, as in the best-fit torque model discussed above. The morphology is similar to that obtained by K08, but the wakes do not have tapered, elongated tails in the simulation. We can also see that $\lambda$ is higher near the particles in the simulation than it is for K08. This makes sense because K08 assume the perturbers to be weak ($\lambda\ll 1$), but in reality they are not (see Section 5.2.2 below).
@@ 4.8 Spatial variation of key quantities and its evolution
The top row shows the mass density in the orbital plane. The white contour is the equidensity surface ($\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}$) that we fit with an ellipse, whereas the grey contours show a few other equidensity surfaces. The softening spheres of the AGB core (left) and companion (right) are represented by black circles. Note that the axis units are normalized to the current particle separation. Hence, the particle separation is always equal to $1$ in these units, but the softening radius, which is constant ($r_{\mathrm{soft}}=2.41\,\mathrm{R_{\odot}}$), appears to grow with time in these normalized units. The left and right snapshots are very similar despite $29$ orbits elapsing between them, which implies a high degree of self-similarity in the problem as the separation decreases with time. Note that the AGB core particle has a somewhat higher concentration of gas around it than the companion particle, which makes sense given that the AGB core particle starts the simulation surrounded by dense gas, whereas the companion particle starts the simulation outside the AGB star.
The second row shows the average Mach number of the perturbers $(\mathcal{M}_{1}+\mathcal{M}_{2})/2$. The gas in the torque-dominating region, delineated by the $\rho=\rho_{\mathrm{c}}$ contour, is subsonic. Note the spiral shock, particularly evident in the middle panel, where a sudden transition from blue (high sound speed) to red (low sound speed) is visible across the shock. A jump in density across the shock can be seen in the middle panel of the top row where density contours almost overlap.
@@ 5 SUCCESSES AND LIMITATIONS
We have seen that the simple models presented can be used to understand the torque evolution in our CE simulation. The modeled torque and that measured from the simulation show a remarkable level of agreement. However, for each model there are parameters – e.g. $\langle\widetilde{B}/\widetilde{A}\rangle$, $\overline{\rho}(t)$, and $\langle\Delta\phi\rangle$ for the E04 model and $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$ for the K08 model – that need to be estimated from the simulation, and these estimates entail choices. Further, both types of model rely on assumptions that drastically simplify the problem; to what extent are these assumptions justified?
@@ 5.1 Ellipsoid model
The simplest version of the ellipsoid model is the constant aspect ratio, constant lag angle prolate spheroid model, akin to the model used by E04. This minimalistic model depends on three parameters, viz. $\widetilde{A}/a$, which is time-dependent, and $\widetilde{B}/\widetilde{A}$ and $\Delta\phi$, which are both treated as constants. Recall that $\widetilde{A}/a$ is relevant only because it affects the value of $\overline{\rho}$ (see the discussion of Newton’s third theorem in Section 3.1). This model does a reasonable job of producing the simulation results, particularly the period-averaged magnitude of the torque, which deviates from the torque measured in the simulation by $<16\%$ in the time domain studied (compare orange and black lines in the top panel of Figure 2). The model also does a good job of reproducing the amplitude and period of the torque variation, though it does a poor job of reproducing the temporal phase. This minimalistic version of the model is successful because the values of the parameters are fairly constant (Figure 5). However, it is not clear whether this is generally true, given the large diversity of common envelope parameter values in nature. Moreover, we have not tried to motivate the values of the parameters from first principles and knowledge of the initial conditions (but a place to start would be the K08 model).
In any case, it is rather remarkable that the values of the parameters reported by E04, $\widetilde{B}/\widetilde{A}=1/2$, $\Delta\phi=22.5^{\circ}$ are close to the time-averaged values we obtained, viz. $\langle\widetilde{B}/\widetilde{A}\rangle=0.654$, and $\langle\Delta\phi\rangle=14.9^{\circ}$. In addition, we obtain $\langle\widetilde{A}/a\rangle=1.47$; E04 do not report this value, though they do find the behaviour to be self-similar, i.e. $\widetilde{A}/a\approx\mathrm{const}$. In fact, when we measure the properties of the ellipse in their figure 8, we find $A/a\approx 1.43$, $B/A=0.658$, and $\Delta\phi=16.1^{\circ}$, values that are remarkably similar (almost identical) to the ones we obtain. This near-perfect correspondence is likely a coincidence, but it could suggest that the parameter values tend to converge to values that are fairly universal, in the sense that they only depend on a few basic quantities, such as $\mathcal{M}_{\mathrm{p}}$.
A correspondence of merger phenomena on vastly different scales is further suggested by the morphology in gas density seen in the merging galaxy simulations of Bortolas et al. (2018); see the pre-merger snapshots shown in their figure 2, which are strikingly similar to those seen in our simulation.
@@ 5.2 Double perturber model
Unlike the ellipsoid model which is partly phenomenological, the K08 model is motivated from first principles but relies on various assumptions, discussed below in Section 5.2.2. Aside from these assumptions, an important limitation is that the parameters of the model, viz. the density and sound speed of the unperturbed medium, are difficult to constrain.
@@ 5.2.1 Model parameters
[…]
@@
The torque is also affected by the resolution, and Gagnier & Pejcha (2024) find that one needs $\gtrsim 47$ resolution elements per spline softening radius for convergence, whereas our simulation has a constant value of $r_{\mathrm{soft}}/\delta_{5}\approx 34$. Less resolution per softening length tends to reduce the drag force (Ohlmann, 2016; Chamandy et al., 2019b; Gagnier & Pejcha, 2024). Evidently, solutions can also depend on the locations, sizes, and shapes of the regions where the different levels of resolution are implemented (for example, one can compare models D and F of Chamandy et al. 2018). In this work, the resolution was chosen partly to ensure a reasonable level of energy conservation, although by simulation end, the total energy had increased by $6\%$ (see C).
@@
In any case, the magnitude of the torque is reproduced if one sets $\rho_{0}=\xi\rho_{\mathrm{c}}$ in the K08 model with $\xi=0.44$. While $\xi$ might approach $1$ at late times, this might only be because the torque in the simulation is overestimated. Importantly, in addition to the overall magnitude, the amplitude and phase of the torque are also reproduced remarkably well, which is encouraging. In summary, one obtains an excellent fit by taking $\mathcal{M}_{\mathrm{p}}$ as the value calculated using the mean sound speed in the torque-dominating region and $\rho_{0}$ as $0.44$ times the value of the density at the boundary of this region.
5.2.2 Consequences of neglecting various effects in the K08 model
@@
The agreement between the K08 model and the simulation is rather remarkable given all of the idealizations in the former. For one, the K08 model neglects gas self-gravity, which is justified in the regime considered since the point masses dominate the potential. It is reasonable to neglect self-gravity if the ratio of the orbital period to the gas free-fall time, $P_{\mathrm{orb}}/t_{\mathrm{ff}}\sim\sqrt{G\overline{\rho}}/\Omega\sim\sqrt{\overline{\rho}a^{3}/m}$ is small. From Figure 6, $\overline{\rho}\lesssim 1.8\times 10^{-5}\,{\rm g\,cm^{-3}}$, and from Figure 1, $a\lesssim 23\,\mathrm{R_{\odot}}$. Thus, we obtain $P_{\mathrm{orb}}/t_{\mathrm{ff}}\lesssim 0.2$ at $t\approx 125\,{\rm d}$ and smaller thereafter, so self-gravity likely plays only a minor role.
@@
Secondly, the K08 model solves a linearized set of equations which depend on the assumption that the density perturbations are small, i.e. the perturbers are weak. This constraint is relaxed and the full nonlinear equations are solved by Kim & Kim (2009), who study a single strong perturber in rectilinear motion, and by Kim (2010), who studies a single strong perturber in circular motion. However, the double perturber case has not yet been studied in the nonlinear regime. According to those papers, to be in the linear regime one requires $\mathcal{A}=GM_{\mathrm{p}}/c_{\mathrm{s,0}}^{2}r_{\mathrm{s}}\ll 1$, where $r_{\mathrm{s}}$ is the Plummer softening radius. In our CE simulation, $\mathcal{A}\approx 12$ (taking $r_{\mathrm{soft}}=2.8r_{\mathrm{s}}$), so we are in the nonlinear regime. However, for $\mathcal{M}_{\mathrm{p}}<1$, which is the regime in which we find ourselves (bottom panel of Figure 2), solutions of the drag force are found not to differ very much from the linear case (Kim & Kim, 2009). [^4] For $\mathcal{M}_{\mathrm{p}}>1$, the torque in the nonlinear solutions is smaller than that in the linear solutions. This likely helps to explain why the K08 model works so well to explain our CE simulation, up to a scaling factor of order unity. Nonlinear effects can also become important as $\mathcal{B}=Gm/c_{\mathrm{s,0}}^{2}a$ increases, but for Mach numbers $>1$ (Kim, 2010). Even if the drag torque in the subsonic regime is well approximated by the linear equations, the morphology of the wakes produced by solving the full nonlinear equations, including weak shocks outside the orbit, matches better with what is obtained in our simulations (c.f. the top two panels of figure 5 of Kim 2010).
@@
A third idealization of K08 is that the unperturbed density and sound speed are spatially constant, which at first glance seems inconsistent with the CE case, where the density scale height of the original AGB profile can be comparable to the inter-particle separation during the simulation (Chamandy et al., 2019a). However, the initial density profile gets drastically modified during the CE phase (e.g. Iaconi et al., 2017; Chamandy et al., 2020), and hence seems fairly irrelevant for the late times explored in this work. There seems to be little memory of the original profiles and flow pattern, aside from their influence on how much gas is present near the perturbers at later times.
@@
Another idealization of K08 is that the CM of the binary is not in motion with respect to the envelope. In our CE simulation, the particle CM moves with a speed of a few $\,{\rm km\,s^{-1}}$ (as can be estimated from the inset of Figure 1; see also Chamandy et al. 2019b). As this speed is small compared to the orbital speed of $\sim 60\,{\rm km\,s^{-1}}$ (Figure 9), the CM motion is not expected to affect significantly the torque on the binary (Sánchez-Salcedo & Chametla, 2014).
@@
In the K08 model the semi-major axis of the binary orbit is fixed, whereas in our CE simulation it is decreasing with time. The timescale $\overline{a/\dot{a}}$, where bar here denotes average over an orbital period, is much larger than the orbital period $P_{\mathrm{orb}}$. The flow is expected to reach a new steady state within a sound-crossing time (Kim & Kim, 2007; Desjacques et al., 2022),
@@
In fact, the parameters ($\rho_{0}$ and $\mathcal{M}_{\mathrm{p}}$) also vary on the shorter timescale $P_{\mathrm{orb}}$ due to the orbital eccentricity (e.g. Szölgyén et al., 2022), but relation (26) shows us that even this timescale is large compare to $t_{\mathrm{s}}$. Hence, to a good approximation, one can treat the flow as quasi-steady and apply the K08 model at any point in time during the time interval studied. This is only possible because the perturbers move subsonically at these times. For recent studies of gaseous dynamical friction on perturbers in eccentric orbits see Buehler et al. (2024) and O’Neill et al. (2024). [^5] Our results are consistent with the finding of Buehler et al. (2024) that for $e=0.3$ and $\mathcal{M}=0.6$, the analytical solution of Desjacques et al. (2022), which assumes $e=0$, accurately reproduces the drag force in their simulations.
5.2.3 Limitations that are common to the simulation and K08 model

### v3

@@ 4.7 Comparison of the two idealized models
As we have shown, the lagging spheroid model of Escala+04 and the double perturber model of Kim+08 both reproduce the torque in the simulation reasonably well. Thus, they would be expected to be consistent with one another. Indeed, Kim+08 applied their model to the Escala+04 simulation, arguing that their model can be used to explain the Escala+04 results if $\mathcal{M}_{\mathrm{p}}\sim 0.6$. The bottom-left panel of Figure 3 is adapted from Kim+08 (their figure 1a). The quantity plotted in colour is $\log(\mathcal{D})=\log[c_{\mathrm{s,0}}^{2}a/Gm)\lambda]$, which is equal to a constant plus the logarithm of the fractional density perturbation, $\lambda=(\rho-\rho_{0})/\rho_{0}$ (with $\lambda\ll 1$ assumed in Kim+08), for $\mathcal{M}_{\mathrm{p}}=0.6$. From the bottom panel of Figure 2 the relevant part of our simulation typically has $\mathcal{M}_{\mathrm{p}}\sim 0.6$ as well. To compare the ellipsoid model with that of Kim+08, we annotate their figure by drawing an ellipse with parameters set equal to the time-averaged values from our simulation: $\langle A/a\rangle=1.47$, $\langle B/A\rangle=0.649$, and $\langle\Delta\phi=14.9^{\circ}\rangle$ (see also Figure 3). The fairly close correspondence helps to explain why the Escala+04 and Kim+08 models produce similar values of the torque. The time-averaged best fit ellipse in the orbital plane from our simulation is seen to align quite well with the structure obtained for $\log(\mathcal{D})$ by Kim+08, though the latter is not precisely elliptical as the wakes form elongated tails.
The bottom-right panel of Figure 3 shows $\log(\mathcal{D})$ from our simulation at $t=188.7$ d, which is the same snapshot plotted in the top two panels. To calculate $\mathcal{D}$, we assume $\rho_{0}=0.44\rho_{\mathrm{c}}=0.44\times 0.006\rho_{\mathrm{max}}(t)$, as in the best-fit torque model discussed above. The morphology is similar to that obtained by Kim+08, but the wakes do not have tapered, elongated tails in the simulation. We can also see that $\lambda$ is higher near the particles in the simulation than it is for Kim+08. This makes sense because Kim+08 assume the perturbers to be weak ($\lambda\ll 1$), but in reality they are not (see Section 5.2.2 below).
@@ 4.8 Spatial variation of key quantities and its evolution
The top row shows the mass density in the orbital plane. The white contour is the equidensity surface ($\rho=\rho_{\mathrm{c}}=0.006\rho_{\mathrm{max}}$) that we fit with an ellipse, whereas the grey contours show a few other equidensity surfaces. The softening spheres of the AGB core (left) and companion (right) are represented by black circles. Note that the axis units are normalized to the current particle separation. Hence, the particle separation is always equal to $1$ in these units, but the softening radius, which is constant ($r_{\mathrm{soft}}=2.41\,\mathrm{R_{\odot}}$), appears to grow with time in these normalized units. The left and right snapshots are very similar despite $29$ orbits elapsing between them, which implies a high degree of self-similarity in the problem as the separation decreases with time. Note that the AGB core particle has a somewhat higher concentration of gas around it than the companion particle, which is reasonable given that the AGB core particle starts the simulation surrounded by dense gas, whereas the companion particle starts the simulation outside the AGB star.
The second row shows the spatially dependent Mach number of the particles, equal to the particle tangential speed in the rest frame of the particle CM divided by the local sound speed $c_{\mathrm{s}}$. The gas in the torque-dominating region, delineated by the $\rho=\rho_{\mathrm{c}}$ contour, is subsonic. Note the spiral shock, particularly evident in the middle panel, where a sudden transition from blue (high sound speed) to red (low sound speed) is visible across the shock. A jump in density across the shock can be seen in the middle panel of the top row where density contours almost overlap.
@@ 5 SUCCESSES AND LIMITATIONS
We have seen that the simple models presented can be used to understand the torque evolution in our CE simulation. The modeled torque and that measured from the simulation show a remarkable level of agreement. However, for each model there are parameters – e.g. $\langle\widetilde{B}/\widetilde{A}\rangle$, $\overline{\rho}(t)$, and $\langle\Delta\phi\rangle$ for the Escala+04 model and $\rho_{0}(t)$ and $c_{\mathrm{s,0}}(t)$ for the Kim+08 model – that need to be estimated from the simulation, and these estimates entail choices. Further, both types of model rely on assumptions that drastically simplify the problem; to what extent are these assumptions justified?
@@ 5.1 Ellipsoid model
The simplest version of the ellipsoid model is the constant aspect ratio, constant lag angle prolate spheroid model, akin to the model used by Escala+04. This minimalistic model depends on three parameters, viz. $\widetilde{A}/a$, which is time-dependent, and $\widetilde{B}/\widetilde{A}$ and $\Delta\phi$, which are both treated as constants. Recall that $\widetilde{A}/a$ is relevant only because it affects the value of $\overline{\rho}$ (see the discussion of Newton’s third theorem in Section 3.1). This model does a reasonable job of producing the simulation results, particularly the period-averaged magnitude of the torque, which deviates from the torque measured in the simulation by $16\%$ in the time domain studied (compare orange and black lines in the top panel of Figure 2). The model also does a good job of reproducing the amplitude and period of the torque variation, though it does a poor job of reproducing the temporal phase. This minimalistic version of the model is successful because the values of the parameters are fairly constant (Figure 5). However, it is not clear whether this is generally true, given the large diversity of common envelope parameter values in nature. Moreover, we have not tried to motivate the values of the parameters from first principles and knowledge of the initial conditions (but a place to start would be the Kim+08 model).
In any case, it is rather remarkable that the values of the parameters reported by Escala+04, viz. $\widetilde{B}/\widetilde{A}=1/2$ and $\Delta\phi=22.5^{\circ}$, are close to the time-averaged values we obtained, viz. $\langle\widetilde{B}/\widetilde{A}\rangle=0.654$, and $\langle\Delta\phi\rangle=14.9^{\circ}$. In addition, we obtain $\langle\widetilde{A}/a\rangle=1.47$; Escala+04 do not report this value, though they do find the behaviour to be self-similar, i.e. $\widetilde{A}/a\approx\mathrm{const}$. In fact, when we measure the properties of the ellipse in their figure 8, we find $A/a\approx 1.43$, $B/A=0.658$, and $\Delta\phi=16.1^{\circ}$, values that are remarkably similar (almost identical) to the ones we obtain. This near-perfect correspondence is likely a coincidence, but it could suggest that the parameter values tend to converge to values that are fairly universal, in the sense that they only depend on a few basic quantities, such as $\mathcal{M}_{\mathrm{p}}$. This possibility should be investigated in future work.
A correspondence of merger phenomena on vastly different scales is further suggested by the morphology in gas density seen in the merging galaxy simulations of Bortolas+18; see the pre-merger snapshots shown in their figure 2, which are strikingly similar to those seen in our simulation.
[…]
@@ 5.2 Double perturber model
Unlike the ellipsoid model which is partly phenomenological, the Kim+08 model is motivated from first principles but relies on various assumptions, discussed below in Section 5.2.2. Aside from these assumptions, an important limitation is that the parameters of the model, viz. the density and sound speed of the unperturbed medium, are difficult to constrain.
@@ 5.2.1 Model parameters
We found in Section 4.6 that using a density equal to that at the boundary of the torque-dominating region and a sound speed equal to the average value inside this region leads to relatively good agreement with the torque in the simulation. However, these choices are somewhat arbitrary. Recall that the magnitude of the torque and amplitude of its variations for $155\,{\rm d}<t<250\,{\rm d}$ are reproduced remarkably well if one multiplies by the scaling factor $\xi=0.44$, but after $t=270\,{\rm d}$, the agreement is good without needing to include a scaling factor. It may be, however, that the factor of $0.44$ is needed in general, and that the torque in our simulation is overestimated for $t\gtrsim 250\,{\rm d}$ because the ratio of separation to softening radius – in the range $a/r_{\mathrm{soft}}\sim 3.5$$4.5$ – is too small at the end of the simulation. Gagnier+Pejcha25 find that the gravitational torque is overestimated if $a\leq 3.6r_{\mathrm{soft}}$ (where we have converted Plummer to spline softening radius by multiplying by $2.8$, consistent with that work), which is comparable to the values at the end of our simulation. Moreover, Chamandy+19b compared two identical global CE simulations similar to the one in this work but differing in the softening radius by a factor of two, with the number of resolution elements per softening length fixed. Consistent with the above suggestion, they found a slightly higher drag force for the model with larger softening radius, resulting in slightly smaller separation (their figure A1; see also figure C2 and Table C1 of Chamandy+19a). As mentioned in Section 4.4, the sudden decrease in the mean density inside the surface $\rho=\rho_{\mathrm{c}}$ at $t\approx 270\,{\rm d}$ may also be a consequence of the softening radius becoming too large with respect to $a$.
@@
The torque is also affected by the resolution, and Gagnier+Pejcha25 find that one needs $\gtrsim 47$ resolution elements per spline softening radius for convergence, whereas our simulation has a constant value of $r_{\mathrm{soft}}/\delta_{5}\approx 34$. Less resolution per softening length tends to reduce the drag force (Ohlmann16; Chamandy+19a; Gagnier+Pejcha25). Evidently, solutions can also depend on the locations, sizes, and shapes of the regions where the different levels of resolution are implemented (for example, one can compare models D and F of Chamandy+18). In this work, the resolution was chosen partly to ensure a reasonable level of energy conservation, although by simulation end, the total energy had increased by $6\%$ (see D).
@@
In any case, the magnitude of the torque is reproduced if one sets $\rho_{0}=\xi\rho_{\mathrm{c}}$ in the Kim+08 model with $\xi=0.44$. While $\xi$ might approach $1$ at late times, this might only be because the torque in the simulation is overestimated. Importantly, in addition to the overall magnitude, the amplitude and phase of the torque are also reproduced remarkably well, which is encouraging. In summary, one obtains an excellent fit by taking $\mathcal{M}_{\mathrm{p}}$ as the value calculated using the mean sound speed in the torque-dominating region and $\rho_{0}$ as $0.44$ times the value of the density at the boundary of this region.
5.2.2 Consequences of neglecting various effects, as in the Kim+08 model
@@
The good agreement between the Kim+08 model and the simulation suggests that several approximations of the former are valid in the region of parameter space simulated. For one, the Kim+08 model neglects gas self-gravity, which is justified in the regime considered since the point masses dominate the potential. It is reasonable to neglect self-gravity if the ratio of the orbital period to the gas free-fall time, $P_{\mathrm{orb}}/t_{\mathrm{ff}}\sim\sqrt{G\overline{\rho}}/\Omega\sim\sqrt{\overline{\rho}a^{3}/m}$ is small. From Figure 6, $\overline{\rho}\lesssim 1.8\times 10^{-5}\,{\rm g\,cm^{-3}}$, and from Figure 1, $a\lesssim 23\,\mathrm{R_{\odot}}$. Thus, we obtain $P_{\mathrm{orb}}/t_{\mathrm{ff}}\lesssim 0.2$ at $t\approx 125\,{\rm d}$ and smaller thereafter, so self-gravity likely plays only a minor role.
@@
Secondly, the Kim+08 model solves a linearized set of equations which depend on the assumption that the density perturbations are small, i.e. the perturbers are weak. This constraint is relaxed and the full nonlinear equations are solved by Kim+Kim09, who study a single strong perturber in rectilinear motion, and by Kim10, who studies a single strong perturber in circular motion. However, the double perturber case has not yet been studied in the nonlinear regime. According to those papers, to be in the linear regime one requires $\mathcal{A}=GM_{\mathrm{p}}/c_{\mathrm{s,0}}^{2}r_{\mathrm{s}}\ll 1$, where $r_{\mathrm{s}}$ is the Plummer softening radius. In our CE simulation, $\mathcal{A}\approx 12$ (taking $r_{\mathrm{soft}}=2.8r_{\mathrm{s}}$), so we are in the nonlinear regime. However, for $\mathcal{M}_{\mathrm{p}}<1$, which is the regime in which we find ourselves (bottom panel of Figure 2), solutions of the drag force are found not to differ very much from the linear case (Kim+Kim09). [^4] For $\mathcal{M}_{\mathrm{p}}>1$, the torque in the nonlinear solutions is smaller than that in the linear solutions. This likely helps to explain why the Kim+08 model works so well to explain our CE simulation, up to a scaling factor of order unity. Nonlinear effects can also become important as $\mathcal{B}=Gm/c_{\mathrm{s,0}}^{2}a$ increases, but for Mach numbers >1>1 (Kim10). Even if the drag torque in the subsonic regime is well approximated by the linear equations, the morphology of the wakes produced by solving the full nonlinear equations, including weak shocks outside the orbit, matches better with what is obtained in our simulations (c.f. the top two panels of figure 5 of Kim10).
@@
A third idealization of Kim+08 is that the unperturbed density and sound speed are spatially constant, which at first glance seems inconsistent with the CE case, where the density scale height of the original AGB profile can be comparable to the inter-particle separation during the simulation (Chamandy+19b). However, the initial density profile gets drastically modified during the CE phase (e.g. Iaconi+17; Chamandy+20), and hence seems fairly irrelevant for the late times explored in this work. There seems to be little memory of the original profiles and flow pattern, aside from their influence on how much gas is present near the perturbers at later times.
@@
Another idealization of Kim+08 is that the CM of the binary is not in motion with respect to the envelope. In our CE simulation, the particle CM moves with a speed of a few $\rm km\,s^{-1}$} (as can be estimated from the inset of Figure 1; see also Chamandy+19a). As this speed is small compared to the orbital speed of $\sim 60\,{\rm km\,s^{-1}}$ (Figure 10), the CM motion is not expected to affect significantly the torque on the binary (Sanchez-salcedo+Chametla14).
@@
In the Kim+08 model the semi-major axis of the binary orbit is fixed, whereas in our CE simulation it is decreasing with time. The timescale $\overline{a/\dot{a}}$, where bar here denotes average over an orbital period, is much larger than the orbital period $P_{\mathrm{orb}}$. The flow is expected to reach a new steady state within a sound-crossing time (Kim+Kim07; Desjacques+22),
@@
In fact, the parameters ($\rho_{0}$ and $\mathcal{M}_{\mathrm{p}}$) also vary on the shorter timescale $P_{\mathrm{orb}}$ due to the orbital eccentricity (e.g. Szolgyen+22), but relation (26) shows us that even this timescale is large compare to $t_{\mathrm{s}}$. Hence, to a good approximation, one can treat the flow as quasi-steady and apply the Kim+08 model at any point in time during the time interval studied. This is only possible because the perturbers move subsonically at these times. For recent studies of gaseous dynamical friction on perturbers in eccentric orbits see Buehler+24 and Oneill+24.555Our results are consistent with the finding of Buehler+24 that for $e=0.3$ and $\mathcal{M}=0.6$, the analytical solution of Desjacques+22, which assumes $e=0$, accurately reproduces the drag force in their simulations.
5.2.3 Limitations that are common to the simulation and Kim+08 model